OpenStudy (anonymous):

help? =) 2x+2y-z=9 3x+3y+z=16 x-2y=-1

OpenStudy (campbell_st):

well the easy this to do is to rewrite the last equation as x = 2y - 1 and then substitute it into the first 2 equations 2(2y - 1) + 2y - z = 9 and 3(2y - 1) + 3y + z = 16 so your 2 equations are 6y - z = 11 9y + z = 19 you now have 2 equations and 2 unknowns. next step is to eliminate another letter, z is the easiest add the equations together 15y = 30 so y = 2 you just now need to backtrack and find z and then x

OpenStudy (anonymous):

thank you!! =)

OpenStudy (across):

For systems bigger than two equations and two unknowns, it's better to use matrices. Do you know linear algebra?

OpenStudy (anonymous):

matrices? yeah

OpenStudy (across):

This system can be represented in matrix form by\[\begin{bmatrix} 2 & 2 & -1\\ 3 & 3 & 1\\ 1 & -2 & 0 \end{bmatrix}\begin{bmatrix} x\\ y\\ z \end{bmatrix}=\begin{bmatrix} 9\\ 16\\ 1 \end{bmatrix}.\]All you need to do is find the largest matrix's inverse, and you must know the rest from there.

OpenStudy (anonymous):

i had trouble starting off the equations but yall helped so much, thanks